Circular No. 8595

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 Q4 (LINEAR) An apparently asteroidal object discovered by LINEAR (discovery ob-

servation below), and posted on the ‘NEO Confirmation Page’, has been found to show a 16′′ tail in p.a. 255^{o} on CCD images taken by J. Lacruz (Madrid, Spain, 0.30-m reflector) on Sept. 3.04–3.05 UT. J. Young (Table Mountain, 0.6-m f/16 Cassegrain reflector) reports that his long CCD exposures taken on Sept. 4.5 show an extension 16′′ long between p.a. 240^{o} and 270^{o} and a very round, condensed coma 4′′ in diameter in very good seeing; his images on Sept. 5.5 show a 6′′ slightly elongated coma (total mag 17.5) with a 40′′ straight, narrow tail in p.a. 265^{o}.

2005 UT  α_{2000}  δ_{2000}  Mag.
Aug. 31.40325  5^{h}14^{m}52.^{s}38  +33^{o}52′04′′.4  19.1

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2005-R24.

T = 2005 Sept.28.882 TT  ω = 51.^{o}377
e = 0.55154  Ω = 10.811 }2000.0
q = 1.73505 AU  i = 17.531
a = 3.86888 AU n^{o} = 0.129517 P = 7.6 years

COMET P/2005 R1 (NEAT) K. J. Lawrence reports the NEAT discovery of a comet, on CCD images

taken with the 1.2-m Schmidt telescope at Palomar (discovery observation below), whose image appears somewhat softer than surrounding stars and has a possible short tail toward the southwest. Young reports that his images taken on Sept. 5.44–5.45 UT show a 6′′ round coma of mag 17.6 with a broad 14′′ tail spanning p.a. 240^{o}–260^{o}.

2005 UT  α_{2000}  δ_{2000}  Mag.
Sept.2.42614  0^{h}37^{m}07.^{s}81  +31^{o}42′49′′.5  18.0

Preliminary elliptical orbital elements from MPEC 2005-R25:

T = 2005 Oct. 8.552 TT  ω = 117.^{o}622
e = 0.64458  Ω = 259.502 }2000.0
q = 2.06554 AU  i = 15.489
a = 5.81161 AU n^{o} = 0.070349 P = 14.0 years
2005 September 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green